# The One RING [Uncaptioned image]: a Robotic Indoor Navigation Generalist

### v1

@@ 4.5 Ablation Studies
Table 6: Collision Penalty. Adding a small collision penalty (0.1) to the reward function results in 50% less collision, forcing the policy to take more conservative paths.
@@ 5 Conclusion
In this paper, we introduce Ring (Robotic Indoor Navigation Generalist), an embodiment-agnostic policy, trained solely in simulation with diverse randomly initialized embodiments at scale (1M embodiments). Ring displays zero-shot generalization capability to various unseen embodiments, maintaining consistent performance across all. Our experimental results demonstrate that Ring achieves state-of-the-art results on novel embodiments, including in some cases improving over embodiment-specific policies. It can be directly deployed to the real-world despite being solely trained in simulation. Finally, Ring is able to dynamically adjust its behavior based on its embodiment and interactions with the environment.
Supplementary Material
@@
  • Details about the real-world evaluation robot platforms and human evaluation setup (App. 6),
  • Data collection for randomized embodiments using expert planners in simulation (App. 7),
  • Full experimental setup (App. 8),
  • Model architecture details (App. 9),
  • A visualization of the random embodiments in our training set, along with the 5 nearest neighbors to each of the real robots (Stretch-RE1, LoCoBot, Unitree A1) (App. 10),
  • Out-of-distribution generalization for different embodiment parameters (App. 11), and
  • Limitations (conditioning on explicit embodiment parameters(App. 12).
@@
  • Real-world qualitative videos of evaluating Ring zero-shot on four different robot platforms, including Stretch RE-1 with our camera setup, Stretch RE-1 with factory camera configuration, LoCoBot, and Unitree GO1,
@@
Figure 8: Robot platforms. We use $3$ different platforms, including Stretch RE-1, LoCoBot, and Unitree GO1 for our real-world evaluations.
@@
6.1 Stretch RE-1, LoCoBot, Unitree GO1
We use Stretch RE-1, LoCoBot, and Unitree GO1 as our robot platforms for real-world evaluations, shown in Fig. 8. For Stretch RE-1, we evaluate two different sensor configurations: the factory configuration and the configuration suggested by SPOC [17]. We summarize the main differences in these platform in Tab. 7. For robot movements, we either implement a Kalman filter or wrap around provided robot APIs to realize low-level controllers for a discrete action space {MoveBase(±20cm), RotateBase(±$6^{\circ}$, ±$30^{\circ}$), Done} across all platforms. It is important to note that during the training stage, we do not use any embodiment configurations from these robots to generate imitation learning data or to initialize RL fine-tuning embodiments.
Stretch RE-1 Stretch RE-1 (Factory) LoCoBot Unitree GO1
Body dimension (cm) 33 $\times$ 34 $\times$ 141 33 $\times$ 34 $\times$ 141 35$\times$35$\times$89 64.5$\times$28$\times$40
Camera model 2$\times$ D455 D435 D435 D435
Camera vertical FoV (degrees) $59^{\circ}$ $69^{\circ}$ $42^{\circ}$ $42^{\circ}$
Camera horizontal FoV (degrees) $90^{\circ}$ $42^{\circ}$ $68^{\circ}$ $68^{\circ}$
Camera height (cm) 140 130 87 28
Camera pitch (degrees) $27^{\circ}$ $30^{\circ}$ $0^{\circ}$ $0^{\circ}$
@@ 6.1 Stretch RE-1, LoCoBot, Unitree GO1
Figure 9: Human Participants for Navigation Assitance. Our five human participants have different height and different camera-holding poses, resulting in different sensory observations (details about human participants in Table 8).
@@ 6.2 Human Evaluations
Human participants. We asked five human participants to use Ring as a navigation assistant and evaluated its performance across a range of different human embodiments. These variations stem from differences in camera-holding poses, participant heights, step sizes, and rotation angles. A summary of these variations across different participants can be found in Tab.8 and Fig.9. Due to these variations, each participant contributes a unique set of evaluation embodiments and sensor configurations.
Figure 10: iOS app for human evaluation. We developed a simple iOS app that enables human participants to text goal, capture an image using iPhone’s back camera, send both to a remote server, and receive the predicted action from our Ring policy.
Figure 11: Real-world test scene for human evaluation. Three different target objects include (Refer to caption Apple, Refer to caption Houseplant, Refer to caption Mug). The red arrow in the bottom left corner is the starting location.
Human evaluation details. We developed a simple iOS app, as shown in Fig. 10, that enables human participants to input the target object as text (e.g., Find a mug), capture an image using the iPhone’s back camera, and send both the text prompt and the captured image to a remote server. Upon receiving the prompt and image, the remote server processes them using our Ring policy to predict action probabilities and samples an action. The predicted action is then sent back to the iPhone and displayed within the app. The prompted action space used is identical to that of our real-world robot.
The human participant navigates by following the suggested action at their own pace and rotation degree, as detailed in Tab. 8. After completing each action, they tap the Predict button again, prompting the app to capture a new image and send it to the server along with the text prompt. This process repeats until the Done action is returned or the 100th step is reached. An episode is considered successful if, before reaching 100 steps, the target object is visible in the latest image and within a 1 meter distance when Ring calls Done.
The layout of the test scene is shown in Fig. 11, illustrating two locations for finding a Mug, three for a Houseplant, and one for an Apple. The participant always begins at the bottom-left corner of the scene.
@@ 8.1 Real-World Benchmarks
All robots are evaluated in a multi-room apartment shown in Fig. 4. Based on the embodiment, the benchmark has different starting locations and objects. Among our target object categories, Apple can be found in the Living room and Kitchen, Bed can only be found in the Bedroom, Sofa and Television can only be found in the Living room, Vase can be found in the Livingroom, Corridor, Office, and Kitchen, Chair can be found in the Office and Kitchen, HousePlant can be found in the Living room, Office, and Kitchen.
  • LoCoBot: Following Phone2Proc [10], use the same five target object categories, including Apple, Bed, Sofa, Television, and Vase, and the three starting poses shown in 4.
@@ 10 Nearest Neighbor Embodiments to Real Robots in our Training Data
Fig. 13 presents a t-SNE visualization of the embodiment parameters $\mathbf{c}_{e}\in\mathbb{R}^{19}$ for 50k samples from the random embodiments in our training set (examples showin in Fig. 14). We also show the corresponding parameters for Stretch, LoCoBot, and Unitree A1 for visualization purposes. Our random embodiments range widely over the space of possible embodiments, with many closely approximating each of the three real robots. Tables 10, 11, and 12 list the five nearest neighbors to each robot in the compressed t-SNE space and their corresponding embodiment parameters. Although the nearest neighbors do not exactly match each robot’s embodiment, they are sufficiently similar across different parameters. This extensive coverage of the embodiment space and proximity to real-world embodiments ensure consistent zero-shot generalization to all three robots.
@@ 11 Generalization to out-of-distribution embodiment parameters
Stretch Stretch (Factory Config) LoCoBot Unitree A1 Average
@@
12 Limitations
We train Ring-Emb-Cond by explicitly providing the embodiment information to the policy. The embodiment parameters are represented as a configuration vector $\mathbf{c}_{e}\in\mathbb{R}^{19}$, with each dimension corresponding to a specific embodiment parameter listed in Table 1. This information is passed as an additional token to the Transformer State Encoder. We use a simple MLP to project $\mathbf{c}_{e}$ to the desired feature dimension ${e\in\mathbb{R}^{1\times 512}}$ before passing it to the encoder.
Table 14 evaluates the 2 versions of the policy on our custom benchmark consisting of $2{,}000$ random embodiments across $2{,}000$ scenes, comparing metrics such as Success Rate, Success Weighted by Collision (SC), Collision Rate (CR), and Safe Episode (percentage of episodes without any collisions).
The results do not show a clear benefit to conditioning the policy on embodiment information. This could be due to several reasons. It is possible that most relevant information about environment hazards and agent motion can be already inferred from visual observations. It is also possible that a significant fraction portion of collisions (both with an without embodiment specification provided) occur with objects that never enter the agent’s visual field, in which case extra information about its own embodiment would not help. Alternatively, a more effective method for conditioning the policy on the parameters may exist. Future work should explore this with additional examination of agent-environment collision and and designing improved policy architectures to better integrate embodiment parameters, ultimately training a more efficient and robust policy that explicitly incorporates embodiment information.
Body Config
Ring 67.62 56.24 42.53 7.77 46.90
Ring-Emb-Cond 69.44 57.42 44.69 8.0 46.54

### v2

@@ 4.4 Ring changes its behavior across different embodiments
Table 6: Collision Penalty. Adding a small collision penalty (0.1) to the reward function results in 50% less collision, forcing the policy to be even more conservative.
4.5 Collision analysis
Ring is trained with randomized body dimensions and is not explicitly provided with real embodiment information. Regardless, the learned policy remains as safe—and in some cases even safer—than embodiment-specific policies (Tab. 6). Collision-avoidance behavior can be further improved by incorporating a small collision penalty into the RL reward (Tab. 6).
Ring learns to take conservative paths without explicit knowledge of its collider size. Expert trajectories are generated using A*, finding minimum-cost paths where the cost is defined as the inverse of the Euclidean distance to the nearest obstacles. During reinforcement learning, collisions slow down the agent’s progress, leading to lower rewards through a step penalty. Since collisions cause no meaningful state changes and only waste time, the policy learns to avoid them to complete tasks more efficiently. We evaluate Ring against embodiment-specific baselines using the metric Safe Episodes—the percentage of episodes completed without any collisions. As shown in Tab. 6, Ring achieves a higher percentage of Safe Episodes compared to the baselines. By training across a large number of embodiments and without access to exact body size information, Ring learns to take more conservative actions through reinforcement learning, promoting safer navigation.
Include collision penalty to take safer routes. Adding a small collision penalty of $0.1$ to the reward function can further reduce collision rate (CR) by $50\%$. The resulting policy is more conservative, regardless of embodiment size. To quantify these results, we created a custom benchmark similar to Chores-$\mathbb{S}$ [12], consisting of $2{,}000$ random embodiments across $2{,}000$ scenes. We evaluate $2$ different versions of our policy on this benchmark, comparing metrics such as Success Rate, Success Weighted by Collision (SC), Collision Rate (CR), and Safe Episode. As shown in Tab. 6, adding the collision penalty reduces the collision rate (CR) ($7.77\%$rightarrow $4.03\%$) as well as increases the percentage of trajectories without collisions ($46.90\%$rightarrow $60.57\%$).
@@ 5 Conclusion
In this paper, we introduce Ring (Robotic Indoor Navigation Generalist), an embodiment-agnostic policy trained entirely in simulation on 1 million diverse, randomly initialized embodiments. Ring demonstrates strong zero-shot generalization to unseen embodiments, maintaining consistent performance across a wide range of robots. It achieves state-of-the-art results on novel embodiments—sometimes even outperforming embodiment-specific policies—and transfers directly to the real world despite never seeing real robot embodiments during training. Finally, Ring dynamically adapts its behavior based on its embodiment and interactions with the environment.
6 Limitations
We train Ring-Emb-Cond by explicitly providing the embodiment information to the policy. The embodiment parameters are represented as a configuration vector $\mathbf{c}_{e}\in\mathbb{R}^{19}$, with each dimension corresponding to a specific embodiment parameter listed in Table 1. This information is passed as an additional token to the Transformer State Encoder. We use a simple MLP to project $\mathbf{c}_{e}$ to the desired feature dimension ${e\in\mathbb{R}^{1\times 512}}$ before passing it to the encoder. Tab. 7 evaluates the 2 versions of the policy on our custom benchmark consisting of $2{,}000$ random embodiments across $2{,}000$ scenes, comparing metrics such as Success Rate, Success Weighted by Collision (SC), Collision Rate (CR), and Safe Episode (percentage of episodes without any collisions).
The results do not show a clear benefit to conditioning the policy on embodiment information. This could be due to several reasons. It is possible that most relevant information about environment hazards and agent motion can be already inferred from visual observations. It is also possible that a significant fraction portion of collisions (both with an without embodiment specification provided) occur with objects that never enter the agent’s visual field, in which case extra information about its own embodiment would not help. Alternatively, a more effective method for conditioning the policy on the parameters may exist. Future work should explore this with additional examination of agent-environment collision and designing improved policy architectures to better integrate embodiment parameters, ultimately training a more efficient and robust policy that explicitly incorporates embodiment information.
Embodiment Parameters
Ring 67.62 56.24 42.53 7.77 46.90
Ring-Emb-Cond 69.44 57.42 44.69 8.0 46.54
Appendices for The One RING [Uncaptioned image]: a Robotic Indoor Navigation Generalist
@@ 6 Limitations
  • Details about the real-world evaluation robot platforms and human evaluation setup (App. 7),
  • Data collection for randomized embodiments using expert planners in simulation (App. 8),
  • Full experimental setup (App. 9),
  • Model architecture details (App. 10),
  • We study the impact of using a more powerful visual encoder (App. 11),
  • A visualization of the random embodiments in our training set, along with the 5 nearest neighbors to each of the real robots (Stretch-RE1, LoCoBot, Unitree A1) (App. 12),
  • Out-of-distribution generalization for different embodiment parameters (App. 13)
@@
  • Real-world qualitative videos of evaluating Ring zero-shot on five different robot platforms, including Stretch RE-1 with our camera setup, Stretch RE-1 with factory camera configuration, LoCoBot, Unitree GO1, and RB-Y1 wheeled humanoid,
@@
7.1 Stretch RE-1, LoCoBot, Unitree GO1, RB-Y1
We use Stretch RE-1, LoCoBot, Unitree GO1, and RB-Y1 wheeled humanoid as our robot platforms for real-world evaluations, shown in Fig. 8. For Stretch RE-1, we evaluate two different sensor configurations: the factory configuration and the configuration suggested by SPOC [12]. For RB-Y1, we include both standing and seated configurations and use an iPhone 16 Pro camera mounted on the robot to stream visual observations. The main differences among these platforms are summarized in Tab. 8. For robot movements, we either implement a Kalman filter or wrap around provided robot APIs to realize low-level controllers for a discrete action space {MoveBase(±20cm), RotateBase(±$6^{\circ}$, ±$30^{\circ}$), Done} across all platforms. It is important to note that during the training stage, we do not use any embodiment configurations from these robots to generate imitation learning data or to initialize RL fine-tuning embodiments.
Figure 7: textbfiOS app for human evaluation. We developed a simple iOS app that enables human participants to text goal, capture an image using iPhone’s back camera, send both to a remote server, and receive the predicted action from our Ring policy.
Stretch RE-1 Stretch RE-1 (Factory) LoCoBot Unitree GO1 RB-Y1 (standing) RB-Y1 (seated)
Body dimension (cm) 33$\times$34$\times$141 33$\times$34$\times$141 35$\times$35$\times$89 64.5$\times$28$\times$40 60$\times$69$\times$140 60$\times$69$\times$92
Camera model 2$\times$ D455 D435 D435 D435 iPhone 16 Pro Camera iPhone 16 Pro Camera
Camera vertical FoV $59^{\circ}$ $69^{\circ}$ $42^{\circ}$ $42^{\circ}$ $73^{\circ}$ $73^{\circ}$
Camera horizontal FoV $90^{\circ}$ $42^{\circ}$ $68^{\circ}$ $68^{\circ}$ $53^{\circ}$ $53^{\circ}$
Camera height (cm) 140 130 87 28 140 92
Camera pitch $27^{\circ}$ $30^{\circ}$ $0^{\circ}$ $0^{\circ}$ $0^{\circ}$ $0^{\circ}$
@@ 7.2 Human Evaluations
Human participants. We asked five human participants to use Ring as a navigation assistant and evaluated its performance across a diverse range of human embodiments. These embodiment variations arose from differences in camera-holding posture, participant height, step size, and rotation angle. Tab. 9 summarizes these variations across participants. As a result, each participant contributed a distinct set of evaluation embodiments and sensor configurations.
Human Evaluation Details. We developed a simple iOS app (Fig. 7) that allows human participants to input a text prompt (e.g., “Find a mug”), capture an image using the iPhone’s back camera, and send both the prompt and image to a remote server. The server processes this input using our Ring policy, predicts action probabilities, samples an action, and returns it to the app for display. The action space available in the app mirrors that of our real-world robot.
Participants follow the suggested action at their own pace and chosen rotation degree, as specified in Tab. 9. After each step, they tap the Predict button to repeat the process: capturing a new image and sending it, along with the original prompt, to the server. This continues until either the Done action is returned or 100 steps have been executed. An episode is considered successful if the target object is visible in the final image, within 1 meter, when Ring issues the Done action.
Fig. 10 shows the layout of the evaluation scene and two sample trajectories, including two locations for a Mug, three for a Houseplant, and one for an Apple. Participants always begin in the bottom-left corner of the scene (results shown in Tab. 4).
Figure 8: Robot platforms. We evaluate on $6$ different platforms: Stretch RE-1 (with two camera configurations), LoCoBot, Unitree GO1, and the RB-Y1 wheeled humanoid, tested in both standing and seated configurations.
Figure 9: RB-Y1 Trajectories. We deploy the policy on the wheeled humanoid in an unstructured kitchen area (layout shown in Fig. 10) to navigate to different objects. We include both seated and standing configurations and use an iPhone 16 Pro camera to stream the visual observations.
@@ 9.1 Real-World Benchmarks
We evaluate 3 of the robots (Stretch-RE1, LoCoBot, Unitree Go1) in a multi-room apartment shown in Fig. 6. Based on the embodiment, the benchmark has different starting locations and objects. Among our target object categories, Apple can be found in the Living room and Kitchen, Bed can only be found in the Bedroom, Sofa and Television can only be found in the Living room, Vase can be found in the Livingroom, Corridor, Office, and Kitchen, Chair can be found in the Office and Kitchen, HousePlant can be found in the Living room, Office, and Kitchen.
We also deploy the policy on the RB-Y1 wheeled humanoid in an unstructured kitchen environment—similar to the human evaluation setting.
  • LoCoBot: Following Phone2Proc [63], use the same five target object categories, including Apple, Bed, Sofa, Television, and Vase, and the three starting poses shown in 6.
@@
  • RB-Y1: The wheeled humanoid navigates to various target objects in a real kitchen area, including a mug, apple, houseplant, and trashcan (2 example trajectories shown in Fig. 9).
@@ 10 Model Architecture Details
11 A More Powerful Pretrained Visual Encoder.
The default vision encoder used in our policies is the pretrained SIGLIP-ViT-B/16. In this section, we examine the impact of using a more powerful visual encoder on Ring’s performance. We train Ring-Large using OpenAI’s ViT-L/14 336px CLIP model [66]. Table 14 compares the results, showing that a stronger visual encoder significantly improves zero-shot performance across all four embodiments (approximately $9\%$ improvement on average). A larger visual encoder is particularly beneficial in our policy, as the visual observations are highly varied due to randomized camera parameters. To ensure fair comparison with the baselines and because ViT-L/14 is more computationally demanding, we chose to use the ViT-B/16 encoder for our main experiments. We will release the training code for the community for those interested in training with the larger visual encoder.
Model Visual Encoder Benchmark Embodiment
Stretch Stretch (Nav) LoCoBot Unitree A1
Ring SIGLIP-ViT-B/16 76.0 74.0 66.5 72.0
Ring-Large ViT-L/14 336px CLIP 83.8 77.7 75.3 79.9
@@ 12 Nearest Neighbor Embodiments to Real Robots in our Training Data
Fig. 12 presents a t-SNE visualization of the embodiment parameters $\mathbf{c}_{e}\in\mathbb{R}^{19}$ for 50k samples from the random embodiments in our training set (examples showin in Fig. 13). We also show the corresponding parameters for Stretch, LoCoBot, and Unitree A1 for visualization purposes. Our random embodiments range widely over the space of possible embodiments, with many closely approximating each of the three real robots. Tables 12, 12, and 13 list the five nearest neighbors to each robot in the compressed t-SNE space and their corresponding embodiment parameters. Although the nearest neighbors do not exactly match each robot’s embodiment, they are sufficiently similar across different parameters. This extensive coverage of the embodiment space and proximity to real-world embodiments ensure consistent zero-shot generalization to all three robots.
@@ 13 Generalization to out-of-distribution embodiment parameters
Parameter Stretch Stretch (Factory Config) LoCoBot Unitree A1 Average